A solution is prepared by dissolving 5.02 g of KNO 3 in 95.0 g of water. Calculate the

following.

1. Mole fraction of the solute and the solvent

2. % mass of the solute

3. Concentration of solution in ppm.


1
Expert's answer
2021-02-22T05:55:36-0500

1. n1=5.02/101 = 0,0497029703 moles.

n2= 95/18 = 5,27777778 moles.

n1 x 100%/n1+n2 = 0,0497029703 x 100/ 0,0497029703 + 5,27777778 = 0,93295448%. of KNO3.

99,0670455 % of H2O.

2. W% = 5.02 x 100% / 5.02 + 95 = 5,0189962 %.

3. 50100 ppm.